# SPECIAL RELATIVITY

### v9

@@ 15.3.5 Experimental Test of SR Theory in 3rd Generation Light Source
One way to demonstrate incompatibility between the standard approach to relativistic electrodynamics, which deals with the usual Maxwell’s equations, and particle trajectories calculated by using non-covariant particle tracking, is to make a direct laboratory test of synchrotron radiation theory. In other words, we are stating here that, despite the many measurements done during decades, synchrotron radiation theory is not an experimentally well-confirmed theory.
Let us analyze the potential for exploiting synchrotron radiation sources in order to confirm the predictions of corrected synchrotron radiation theory. The emittance of the electron beam in new generation synchrotron radiation sources is small enough, so that one can neglect finite electron beam size and angular divergence in the soft X-ray wavelength range, and such synchrotron radiation source can be examined under the approximation of a filament electron beam. This allows us to take advantage of analytical presentations for single electron synchrotron radiation fields.
@@
The basic setup for a test experiment is sketched in Fig 67. The soft X-ray undulator beam line should be tuned to a minimum photon energy (typically this limit is related with the so called ”water window” wavelength range). The radiation pulse goes through a monochromator filter $F$ and its energy is subsequently measured by the detector. No precise monochromatization of the undulator radiation is required in this case: a monochromator line width $\Delta\omega/\omega\simeq 10^{-3}$ is sufficient. In order for proposed test experiment to be curried out, it is necessary to control the beam kicking e.g. by corrector magnet. In the case of no kick the maximum pulse energy registered by the detector will coincide with the monochromator line tuned to resonance. When the kick is introduced the conventional synchrotron radiation theory still predicts a zero red shift in the resonance wavelength. In contrast to this, one of the immediate consequences of the corrected theory is the occurrence of a non-zero red shift.
The proposed experimental procedure is relatively simple, because is based on relative measurements in the (electron beam) velocity direction with and without transverse kick. Such a measurement is critical, in the sense that the prediction of conventional theory is the absence of red shift and (to our knowledge) has never been performed at synchrotron radiation facilities. However, XFEL based experiment confirm our correction for spontaneous undulator emission (1).
@@ 15.4 Synchrotron Radiation from Bending Magnets
Consider a single relativistic electron moving on a circular orbit. The observer in the standard treatment is assumed to be located in a vertical plane tangent to the circular trajectory at the origin, at an angle $\theta$ above the level of the orbit. In other words, in this geometry the $z$ axis is not fixed, but depends on the observer’s position. Note that the geometry of the electron motion has a cylindrical symmetry, with the vertical axis going through the center of the circular orbit. Because of this symmetry, in order to calculate spectral and angular photon distributions, it is not necessary to consider an observer at a more general location.
There are a number of remarkable effects which are a consequence of the cylindrical symmetry and the paraxial approximation. In discussing this we will demonstrate that there is no needs to use covariant particle tracking for derivation of the bending magnet radiation. However, there is one situation where the conventional theory fails. The covariant approach predicts a non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion along the magnetic field (i.e with respect to nominal orbit).
It would be well to begin with bird’s view of the main results. We want to solve the electrodynamics problem based on Maxwell’s equations in their usual form. In this case we should analyze the particle evolution within the framework of special relativity, where the problem of assigning Lorentz coordinates to the lab frame in the case of accelerating motion is complicated. The only possibility to introduce Lorentz coordinates in this situation consists in introducing individual coordinate systems (i.e. individual rule-clock structure) for each point of the path.
In order to keep Lorentz coordinates in the lab frame, as discussed before, we need only to perform a clock resynchronization by introducing a time shift and change the scale of time, that is the rhythm of clocks. Because of cylindrical symmetry, the observer is assumed to be located in the vertical plane tangent to the circular trajectory at the origin. Because the formation length is very short, it is necessary to know the velocity and position over only a small arc of the trajectory. In the ultrarelativistic (paraxial) approximation, the observer sees a uniform acceleration of the electron $a=v^{2}/R\simeq c^{2}/R$ in the transverse direction. We can, then, write velocity and offset of the electron as follows $v_{\perp}=at$, $r_{\perp}=at^{2}/2$. Relativity of simultaneity gives time shift $\Delta t=v_{\perp}r_{\perp}/c^{2}=v_{\perp}^{2}t/(2c^{2})$. But a correction of the rhythm is $t\sqrt{1-v_{\perp}^{2}/c^{2}}\simeq t-v_{\perp}^{2}t/(2c^{2})$. We have a beautiful cancellation. This is a coincidence. It is because we have deal with uniform acceleration in the transverse direction using a second order (paraxial) approximation when an electron is moving along an arc of a circle.
@@ 15.4.1 Conventional Theory
Consider a single relativistic electron moving on a circular orbit and an observer. It is worth to underline the difference between the more general geometry which we use and the geometry used in most synchrotron radiation textbooks for the treatment of bending magnet radiation. In standard treatments of bending magnet radiation, the horizontal observation angle $\theta_{x}$ is always equal to zero. The reason for this is that most textbooks focus on the calculation of the intensity radiated by a single electron in the far zone, which involves the square modulus of the field amplitude, but do not analyze, for instance, situations like source imaging.
@@
In standard treatments of bending magnet radiation, the phase term $\exp(i\Phi_{0})$ is absent. In fact, the horizontal observation angle $\theta_{x}$ is always equal to zero.
@@ 15.4.2 Why did the Error in Synchrotron Radiation Remain so Long Undetected?
Our case of interest is an ultrarelativistic electron accelerating in a circle. As already remarked, in conventional (non-covariant) particle tracking the description of the dynamical evolution in the lab frame is based on the use of the absolute time convention. In this case simultaneity is absolute, and we only need one set of synchronized clocks in the lab frame, to be used for the description of the accelerated motion. However, the use of the absolute time convention automatically implies the use of much more complicated field equations, and these equations are different for each value of the particle velocity i.e. for each point along its path. This is the reason to prefer the covariant approach within the framework of both dynamics and electrodynamics.
@@
We start by considering an electron moving along a circular trajectory that lies in the $(x,z)$-plane and tangent to the $z$ axis. Because of cylindrical symmetry, in order to calculate spectral and angular photon distributions, it is not necessary to consider an observer at general location. The observer is assumed to be located in the vertical plane tangent to the circular trajectory at the origin. In ultrarelativistic (paraxial) approximation we evaluate transformations working only up to the order of $v^{2}_{x}/c^{2}$. The restriction to this order provides an essential simplicity of calculations. We can interpret manipulation with rule-clock structure in the lab frame simply as a change of variables according to the transformation Eq. (13): $x_{L}=\gamma_{x}x$, $t_{L}=(t/\gamma_{x}+\gamma_{x}xv_{x}/c^{2})$. We are dealing with a second order approximation and $\gamma_{x}=1+v_{x}^{2}/(2c^{2})$. The overall combination of Galilean transformation and variable changes actually yields to the transverse Lorentz transformation. Since the Galilean transformation, completed by the introduction of the new variables, is mathematically equivalent to a Lorentz transformation, it obviously follows that transforming to new variables leads to the usual Maxwell’s equations.
@@
In order to keep Lorentz coordinates in the lab frame, as discussed before, we need only to perform a clock resynchronization by introducing the time shift $\Delta t=t_{L}-t=-[v_{x}^{2}/(2c^{2})]t+xv_{x}/c^{2}$. The relativistic correction to the particle’s offset ”$x$” does not appear in this expansion order, but only in order of $v_{x}^{3}/c^{3}$ and $x_{L}=x$ in our case of interest. Although we have only shown that time shift in one rather special case, the result is right for any offset and (transverse) velocity direction: $\Delta t=t_{L}-t=-[|\vec{v}_{\perp}|^{2}/(2c^{2})]z^{\prime}/c+\vec{r}_{\perp}\cdot\vec{v}_{\perp}/c^{2}$. To finish our analysis we need only find a relativistic correction to the longitudinal motion. We remark again that if we evaluate the transformations up to the second order $(v_{\perp}/c)^{2}$, the relativistic correction in the longitudinal motion does not appear in this approximation. We have demonstrated the covariant method that can be used for any trajectory - a general way of funding what happens directly in space-frequency domain and in paraxial approximation.
@@
Let us now see how to apply this covariant method to a special situation. Let’s use our knowledge of the relativistically correct method for calculating synchrotron radiation emission to find the photon angular-spectral density distributions from a bending magnet. In the ultrarelativistic approximation, we have a uniform acceleration of the electron $a=v^{2}/R\simeq c^{2}/R$ in the transverse direction. We can, then, write velocity and offset of the electron as follows $v_{x}=at=az^{\prime}/v=az^{\prime}/c$, $x=at^{2}/2=az^{\prime 2}/(2c^{2})$. We have now all quantities we wanted. Let us put them all together in relativistic time shift: $\Delta t=t_{L}-t=-a^{2}z^{\prime 3}/(2c^{5})+a^{2}z^{\prime 3}/(2c^{5})=0$. There is no difference! We do not need to use covariant particle tracking for derivation of the bending magnet radiation. Usually, such a beautiful cancellation is found to stem from a deep underlying principle. Nevertheless, in this case there does not appear to be any such profound implication. This is a coincidence.
@@
This cancellation is not surprising, if one analyzes the general expression for the radiation field from bending magnet in the far zone Eq.(78). In our previous discussion of undulator radiation, we learned that the relativistic correction appears only when the transverse electron trajectory is included in the total phase $\Phi_{T}$ under the integral Eq.(38). Referring back to Eq.(40) for the phase factor $\Phi_{T}$ , we see that the term which depends on the transverse position of the electron can be written as $\exp i(\omega/c)[\theta_{x}x(z^{\prime})+\theta_{y}y(z^{\prime})]$. We conclude that the observation angle in the total phase factor under the integral must be related with the contribution of the transverse electron trajectory. Now look at Eq.(78). This equation includes only the observation angle $\theta_{y}$ in the phase factor under the integral. This means that the transverse constraint motion of the electron in the bending magnet does not affect synchrotron radiation. So we are justified using a non-covariant approach for considering the constrained electron motion along the nominal orbit in $(x,z)$-plane.
@@
We point out that the cancellation in relativistic time shift and the independence of the Fraunhofer propagator (to be more precise, in space-frequency domain we are dealing with a paraxial approximation of Green’s function of nonhomogeneous Helmholtz equation) on the observation angle $\theta_{x}$ in the far zone can be regarded as the two sides of the same coin: they are manifestation of the cylindrical symmetry when an electron is moving along an arc of a circle. Because of cylindrical symmetry, in order to calculate spectral and angular photon distributions in the far zone, it is not necessary to consider an observer at a general location. The observer is assumed to be located in the vertical plane tangent to the circular trajectory at the origin. In this case observation angle $\theta_{x}=0$ and the observation angle $\theta_{y}$ is above the level of the orbit. In other words, in this very special geometry the $z$-axis is not fixed, but depends on the observer position. However, this way of proceeding can hardly help to obtain radiation fields in the near zone. Indeed, in the near zone we are dealing with the Fresnel propagator, which obviously depends on the constrained motion of the electron. We use far-zone arguments only to show that there is no influence of the difference between the non-covariant and covariant trajectories on the synchrotron radiation from bending magnets. The cancellation in the relativistic time shift leads to the same outcome in the near zone as it must be.
@@ 15.4.3 Influence of the Kick According to Conventional Theory
Up to this point we considered an electron moving along a circular trajectory that lies in the $(x,z)$-plane and tangent to the $z$ axis. The phase difference in the fields will be determined by the position of the observer and by the electron trajectory. Let us now discuss the bending magnet radiation from a single electron with arbitrary angular deflection and offset with respect to the nominal orbit.
Approximation for the electron path (see Eq. (168), Eq. (171) in the Appendix III) can be used to characterize the field from an electron moving on any trajectory. Using Eq. (70) and Eq. (71) an approximated expression for $s(z)$ can be found:
@@
According to the conventional (incorrect) coupling of fields and particles, there is a prediction of radiation theory concerning to the bending magnet radiation from a single electron with and without kick. Namely, when a kick is introduced, there is a rigid rotation of the angular distribution in the far zone.
@@ 15.4.4 Influence of the Kick According to Correct Coupling of Fields and Particles
We would like to make a historical note. The difference between covariant and non-covariant particle trajectories was never understood. So, accelerator physicists did not appreciate that there was a contribution to the synchrotron radiation from relativistic kinematics effects. The question now arises how can storage rings actually operate. The point is that this example deals with a situation where electron beam kinetics is determined by the emission of synchrotron radiation from bending magnets. However, because of the cylindrical symmetry, covariant and non-covariant solutions for the electron motion along an arc of a circle yield similar properties of synchrotron radiation except the following modification. The covariant approach predicts a non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion in the vertical direction. But synchrotron radiation from bending magnets is emitted within a wide range of frequencies, and the output intensity is not sensitive on the red shift.

### v10

@@ 19.3.5 Experimental Test of SR Theory in 3rd Generation Light Source
One way to demonstrate incompatibility between the standard approach to relativistic electrodynamics, which deals with the usual Maxwell’s equations, and particle trajectories calculated by using non-covariant particle tracking, is to make a direct laboratory test of synchrotron radiation theory. In other words, we are stating here that, despite the many measurements done over decades, synchrotron radiation theory is not an experimentally well-confirmed theory.
Let us analyze the potential for exploiting synchrotron radiation sources in order to confirm the predictions of corrected synchrotron radiation theory. The emittance of the electron beam in new-generation synchrotron radiation sources is small enough, so that one can neglect finite electron beam size and angular divergence in the soft X-ray wavelength range, and such synchrotron radiation source can be examined under the approximation of a filament electron beam. This allows us to take advantage of analytical presentations for single electron synchrotron radiation fields.
@@
The basic setup for a test experiment is sketched in Fig 56. The soft X-ray undulator beamline should be tuned to a minimum photon energy (typically this limit is related to the so-called ”water window” wavelength range). The radiation pulse goes through a monochromator filter $F$ and its energy is subsequently measured by the detector. No precise monochromatization of the undulator radiation is required in this case: a monochromator line width $\Delta\omega/\omega\simeq 10^{-3}$ is sufficient. For the proposed test experiment to be curried out, it is necessary to control the beam kicking e.g. by corrector magnet. In the case of no kick, the maximum pulse energy registered by the detector will coincide with the monochromator line tuned to resonance. When the kick is introduced the conventional synchrotron radiation theory still predicts a zero red shift in the resonance wavelength. In contrast to this, one of the immediate consequences of the corrected theory is the occurrence of a non-zero redshift.
The proposed experimental procedure is relatively simple because is based on relative measurements in the (electron beam) velocity direction with and without transverse kick. Such a measurement is critical, in the sense that the prediction of conventional theory is the absence of redshift and (to our knowledge) has never been performed at synchrotron radiation facilities. However, XFEL based experiment confirms our correction for spontaneous undulator emission, see the next chapter for more details .
@@ 19.4 Synchrotron Radiation from Bending Magnets
Consider a single relativistic electron moving on a circular orbit. The observer in the standard treatment is assumed to be located in a vertical plane tangent to the circular trajectory at the origin, at an angle $\theta$ above the level of the orbit. In other words, in this geometry, the $z$ axis is not fixed but depends on the observer’s position. Note that the geometry of the electron motion has a cylindrical symmetry, with the vertical axis going through the center of the circular orbit. Because of this symmetry, in order to calculate spectral and angular photon distributions, it is not necessary to consider an observer at a more general location.
There are a number of remarkable effects which are a consequence of the cylindrical symmetry and the paraxial approximation. In discussing this we will demonstrate that there is no need to use covariant particle tracking for derivation of the bending magnet radiation. However, there is one situation where the conventional theory fails. The covariant approach predicts a non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion along the magnetic field (i.e. with respect to nominal orbit).
It would be well to begin with a bird’s view of the main results. We want to solve the electrodynamics problem based on Maxwell’s equations in their usual form. In this case, we should analyze the particle evolution within the framework of special relativity, where the problem of assigning Lorentz coordinates to the lab frame in the case of accelerating motion is complicated. The only possibility to introduce Lorentz coordinates in this situation consists in introducing individual coordinate systems (i.e. individual rule-clock structure) for each point of the path.
In order to keep Lorentz coordinates in the lab frame, as discussed before, we need only to perform a clock resynchronization by introducing a time shift and changing the scale of time, that is the rhythm of clocks. Because of cylindrical symmetry, the observer is assumed to be located in the vertical plane tangent to the circular trajectory at the origin. Because the formation length is very short, it is necessary to know the velocity and position over only a small arc of the trajectory. In the ultrarelativistic (paraxial) approximation, the observer sees a uniform acceleration of the electron $a=v^{2}/R\simeq c^{2}/R$ in the transverse direction. We can, then, write the velocity and offset of the electron as follows $v_{\perp}=at$, $r_{\perp}=at^{2}/2$. Relativity of simultaneity gives time shift $\Delta t=v_{\perp}r_{\perp}/c^{2}=v_{\perp}^{2}t/(2c^{2})$. But a correction of the rhythm is $t\sqrt{1-v_{\perp}^{2}/c^{2}}\simeq t-v_{\perp}^{2}t/(2c^{2})$. We have a beautiful cancellation. This is a coincidence. It is because we have to deal with uniform acceleration in the transverse direction using a second-order (paraxial) approximation when an electron is moving along an arc of a circle.
@@ 19.4.1 Conventional Theory
Consider a single relativistic electron moving on a circular orbit and an observer. It is worth underlining the difference between the more general geometry that we use and the geometry used in most synchrotron radiation textbooks for the treatment of bending magnet radiation. In standard treatments of bending magnet radiation, the horizontal observation angle $\theta_{x}$ is always equal to zero. The reason for this is that most textbooks focus on the calculation of the intensity radiated by a single electron in the far zone, which involves the square modulus of the field amplitude, but do not analyze, for instance, situations like source imaging.
@@
In standard treatments of bending magnet radiation, the phase term $\exp(i\Phi_{0})$ is absent. The horizontal observation angle $\theta_{x}$ is always equal to zero.
@@ 19.4.2 Why did the Error in Synchrotron Radiation Remain so Long Undetected?
Our case of interest is an ultrarelativistic electron accelerating in a circle. As already remarked, in conventional (non-covariant) particle tracking the description of the dynamical evolution in the lab frame is based on the use of the absolute time convention. In this case, simultaneity is absolute, and we only need one set of synchronized clocks in the lab frame, to be used for the description of the accelerated motion. However, the use of the absolute time convention automatically implies the use of much more complicated field equations, and these equations are different for each value of the particle velocity i.e. for each point along its path. This is the reason to prefer the covariant approach within the framework of both dynamics and electrodynamics.
@@
We start by considering an electron moving along a circular trajectory that lies in the $(x,z)$-plane and tangent to the $z$ axis. Because of cylindrical symmetry, in order to calculate spectral and angular photon distributions, it is not necessary to consider an observer at the general location. The observer is assumed to be located in the vertical plane tangent to the circular trajectory at the origin. In ultrarelativistic (paraxial) approximation we evaluate transformations working only up to the order of $v^{2}_{x}/c^{2}$. The restriction to this order provides an essential simplicity of calculations. We can interpret manipulation with rule-clock structure in the lab frame simply as a change of variables according to the transformation Eq. (13): $x_{L}=\gamma_{x}x$, $t_{L}=(t/\gamma_{x}+\gamma_{x}xv_{x}/c^{2})$. We are dealing with a second-order approximation and $\gamma_{x}=1+v_{x}^{2}/(2c^{2})$. The overall combination of Galilean transformation and variable changes actually yields to the transverse Lorentz transformation. Since the Galilean transformation, completed by the introduction of the new variables, is mathematically equivalent to a Lorentz transformation, it follows that transforming to new variables leads to the Maxwell’s equations.
@@
To keep Lorentz coordinates in the lab frame, as discussed before, we need only to perform a clock resynchronization by introducing the time shift $\Delta t=t_{L}-t=-[v_{x}^{2}/(2c^{2})]t+xv_{x}/c^{2}$. The relativistic correction to the particle’s offset ”$x$” does not appear in this expansion order, but only in order of $v_{x}^{3}/c^{3}$ and $x_{L}=x$ in our case of interest. Although we have only shown time shift in one rather special case, the result is right for any transverse velocity direction: $\Delta t=t_{L}-t=-[|\vec{v}_{\perp}|^{2}/(2c^{2})]z^{\prime}/c+\vec{r}_{\perp}\cdot\vec{v}_{\perp}/c^{2}$. To finish our analysis we need only find a relativistic correction to the longitudinal motion. We remark again that if we evaluate the transformations up to the second order $(v_{\perp}/c)^{2}$, the relativistic correction in the longitudinal motion does not appear in this approximation. We have demonstrated the covariant method that can be used for any trajectory - a general way of funding what happens directly in the space-frequency domain and in paraxial approximation.
@@
Let us now see how to apply this covariant method to a special situation. Let’s use our knowledge of the relativistically correct method for calculating synchrotron radiation emission to find the photon angular-spectral density distributions from a bending magnet. In the ultrarelativistic approximation, we have a uniform acceleration of the electron $a=v^{2}/R\simeq c^{2}/R$ in the transverse direction. We can, then, write the velocity and offset of the electron as follows $v_{x}=at=az^{\prime}/v=az^{\prime}/c$, $x=at^{2}/2=az^{\prime 2}/(2c^{2})$. We have now all the quantities we wanted. Let us put them all together in a relativistic time shift: $\Delta t=t_{L}-t=-a^{2}z^{\prime 3}/(2c^{5})+a^{2}z^{\prime 3}/(2c^{5})=0$. There is no difference! We do not need to use covariant particle tracking to derive the bending magnet radiation. Usually, such a beautiful cancellation is found to stem from a deep underlying principle. Nevertheless, in this case, there does not appear to be any such profound implication. This is a coincidence.
@@
This cancellation is not surprising, if one analyzes the general expression for the radiation field from the bending magnet in the far zone Eq.(83). In our previous discussion of undulator radiation, we learned that the relativistic correction appears only when the transverse electron trajectory is included in the total phase $\Phi_{T}$ under the integral Eq.(43). Referring back to Eq.(45) for the phase factor $\Phi_{T}$, we see that the term which depends on the transverse position of the electron can be written as $\exp i(\omega/c)[\theta_{x}x(z^{\prime})+\theta_{y}y(z^{\prime})]$. We conclude that the observation angle in the total phase factor under the integral must be related to the contribution of the transverse electron trajectory. Now look at Eq.(83). This equation includes only the observation angle $\theta_{y}$ in the phase factor under the integral. This means that the transverse constraint motion of the electron in the bending magnet does not affect synchrotron radiation. We are justified in using a non-covariant approach to considering the constrained electron motion along the nominal orbit in the $(x,z)$-plane.
@@
We point out that the cancellation in relativistic time shift and the independence of the Fraunhofer propagator (to be more precise, in the space-frequency domain we are dealing with a paraxial approximation of Green’s function of nonhomogeneous Helmholtz equation) on the observation angle $\theta_{x}$ in the far zone can be regarded as the two sides of the same coin: they are manifestation of the cylindrical symmetry when an electron is moving along an arc of a circle. Because of cylindrical symmetry, in order to calculate spectral and angular photon distributions in the far zone, it is not necessary to consider an observer at a general location. The observer is assumed to be located in the vertical plane tangent to the circular trajectory at the origin. In this case observation angle $\theta_{x}=0$ and the observation angle $\theta_{y}$ is above the level of the orbit. In other words, in this very special geometry, the $z$-axis is not fixed but depends on the observer’s position. However, this way of proceeding can hardly help to obtain radiation fields in the near zone. Indeed, in the near zone, we are dealing with the Fresnel propagator, which depends on the constrained motion of the electron. We use far-zone arguments only to show that there is no influence of the difference between the non-covariant and covariant trajectories on the synchrotron radiation from bending magnets. The cancellation in the relativistic time shift leads to the same outcome in the near zone as it must be.
@@ 19.4.3 Influence of the Kick According to Conventional Theory
Up to this point, we considered an electron moving along a circular trajectory that lies in the $(x,z)$-plane and tangent to the $z$ axis. The phase difference in the fields will be determined by the position of the observer and by the electron trajectory. Let us now discuss the bending magnet radiation from a single electron with arbitrary angular deflection and offset with respect to the nominal orbit.
Approximation for the electron path (see Eq. (155), Eq. (158) in Appendix A3) can be used to characterize the field from an electron moving on any trajectory. Using Eq. (75) and Eq. (76) an approximated expression for $s(z)$ can be found:
@@
According to the conventional (incorrect) coupling of fields and particles, there is a prediction of radiation theory concerning the bending magnet radiation from a single electron with and without the kick. Namely, when a kick is introduced, there is a rigid rotation of the angular distribution in the far zone.
@@ 19.4.4 Influence of the Kick According to Correct Coupling of Fields and Particles
We would like to make a historical note. The difference between covariant and non-covariant particle trajectories was never understood. So, accelerator physicists did not appreciate that there was a contribution to the synchrotron radiation from relativistic kinematics effects. The question now arises how can storage rings operate? The point is that this example deals with a situation where electron beam kinetics is determined by the emission of synchrotron radiation from bending magnets. However, because of the cylindrical symmetry, covariant and non-covariant solutions for the electron motion along an arc of a circle yield similar properties of synchrotron radiation except for the following modification. The covariant approach predicts a non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion in the vertical direction. However synchrotron radiation from bending magnets is emitted within a wide range of frequencies, and the output intensity is not sensitive to the redshift.